OpenStudy (anonymous):

What is the surface area of the right rectangular prism?(Points : 1) 8 m2 12 m2 24 m2 38 m2

OpenStudy (31356):

1. The surface area of the right rectangular prism, S =2(LB+LH+BH) Where L= length, B= breadth and H = height When L=4m ,B=3m, H=1m Surface area oS =2(4×3+4×1+3×1)= 38 m²

OpenStudy (31356):

Medal?

OpenStudy (31356):

Thanks for the medal!

OpenStudy (anonymous):

welcome whaat is ....... What is the volume of the right rectangular prism? (Points : 1) 8 m3 9 m3 12 m3 19 m3

OpenStudy (31356):

Volume of a rectangular prism will be the height x width x length so 4 x 3 x 0.5 = 6m^3

OpenStudy (anonymous):

mmh

OpenStudy (31356):

What is the length, width, and height? That was just an example.

OpenStudy (anonymous):

1m 3m 4m

OpenStudy (31356):

12m3

OpenStudy (31356):

Multiply 1*3*4=12m3 @cayennemocha